What is the pressure of the gas (in atm) when 5.0 moles of CO gas are present in a container of 20.0 L at 27 oC?

A) 3.25 B) 6.15 C) 7.40 D) 9.30 E) 10.55

Please show how you solved.
Answer #1

According to ideal gas law

PV=nRT

P is the pressure

V is the volume

n is number of moles

R is universal gas constant

T is the temperature in kelvin

n= 5 moles

V=20 L

T=27+273 = 300 K

P=(nRT) /V

P=(5*0.0821 L atm mol-1 K-1*300K) /20 L

P=6.15 atm